# Video Event Reasoning and Prediction by Fusing World Knowledge from LLMs with Vision Foundation Models

### v1

@@ 5.6 Discussion and Limitations
The collective results present strong evidence that our framework represents a significant step forward in video understanding. […] Despite these promising results, we acknowledge several limitations that chart clear directions for future research. A primary consideration is the substantial computational cost associated with training and inference, a common challenge for foundation models. Future work should explore advanced model compression techniques (7) and more efficient architectures to democratize access to these capabilities. Moreover, our model’s advanced reasoning is partly derived from our synthetically generated instruction dataset, creating a dependency on the teacher model and a potential for inheriting its latent biases. Developing methods for fine-tuning with more diverse, human-curated data or through direct reinforcement learning from human feedback would be a valuable next step. The persistent issue of factual grounding and hallucination, while mitigated, also warrants further attention. As our qualitative analysis shows, ensuring every statement is perfectly grounded in visual evidence remains an open problem, suggesting a need for tighter fusion mechanisms or post-hoc verification modules. Finally, a notable gap remains between current benchmarks and the unstructured, long-form complexity of real-world video. Scaling our methods to handle hour-long inputs and reason about ambiguous social dynamics, such as those found in datasets like Ego-Exo4D (11), will be a major and exciting challenge for the field.

### v2

@@ 5.5 Discussion and Limitations
The collective results present strong evidence that our framework represents a significant step forward in video understanding. […] Despite these promising results, we acknowledge several limitations that chart clear directions for future research. A primary consideration is the substantial computational cost associated with training and inference, a common challenge for foundation models. Future work should explore advanced model compression techniques and more efficient architectures to democratize access to these capabilities. Moreover, our model’s advanced reasoning is partly derived from our synthetically generated instruction dataset, creating a dependency on the teacher model and a potential for inheriting its latent biases. Developing methods for fine-tuning with more diverse, human-curated data or through direct reinforcement learning from human feedback would be a valuable next step. The persistent issue of factual grounding and hallucination, while mitigated, also warrants further attention. As our qualitative analysis shows, ensuring every statement is perfectly grounded in visual evidence remains an open problem, suggesting a need for tighter fusion mechanisms or post-hoc verification modules. Finally, a notable gap remains between current benchmarks and the unstructured, long-form complexity of real-world video. Scaling our methods to handle hour-long inputs and reason about ambiguous social dynamics, such as those found in datasets like Ego-Exo4D (gandhi2023egoexo4d), will be a major and exciting challenge for the field.